Problem 1: You are considering investing in a 10-year bond issued by NewEnergy Inc. This bond has $1000 face value, 4% coupon rate.
The bond pays coupons semi-annually and is currently selling at $920. The bond can be called at a $1,040 in 3 years.
1.d. (6 points) If the bond is called back by the firm at price of $1,040 in three years, what is the yield to call?
N=
FV=
PMT=
PV=
PMT Type=
Periodic Discount Rate=
Yield to Call =
Answer #1

Using excel formula

Problem 1: You are considering investing in a 10-year bond issued by NewEnergy Inc. This bond has $1000 face value, 4% coupon rate.
The bond pays coupons semi-annually and is currently selling at $920. The bond can be called at a $1,040 in 3 years.
1.d. (6 points) If the bond is called back by the firm at price of $1,040 in three years, what is the yield to call?
N= 20 (=10*20
FV= 1000
PMT= 20 (=1000*4%/2)
PV= 920
PMT Type= 0 (End of Period)
Periodic Discount Rate= 2.51% RATE(20,20,-920,1000)
Yield to Call = 5.35% 2*RATE(20,20,-920,1040)